# Neural Networks for Fast Optimisation in Model Predictive Control: A Review

### v2

@@ Abstract
[iisri]organization=Institute for Intelligent Systems Research and Innovation, Deakin University, addressline=75 Pigdons Rd, city=Waurn Ponds, postcode=3216, state=Victoria, country=Australia
@@ 1 Introduction
Model Predictive Control (MPC) is an optimal control scheme that is widely used in the process industry and robotics. Compared to other control schemes, its main advantages include seamless handling of MIMO systems, ease of consideration of state, input and output constraints, the guarantee of closed loop stability by design, and the ability to exploit knowledge and predictions of future control references or disturbances Schwenzer et al. (2021); Kouvaritakis and Cannon (2015). In a nutshell, MPC controllers use a model or “plant” of the controlled system to derive a sequence of control inputs that will result in optimal system response over a future period of time, or “horizon”, given the current states of the system and desired reference trajectories. This is achieved by solving an optimisation problem where the design variables are the sequence of future system inputs and the plant is used to calculate future system outputs and states for said sequence. The primary challenge in deploying MPC controllers in practical applications stems from the necessity to solve large optimisation problems in real time. For systems with fast dynamics requiring control at high frequency, this control scheme becomes infeasible even for simple systems, especially when the more complex formulations of MPC such as Nonlinear MPC (NMPC) or Robust MPC (RMPC) are used, or when the controller needs to be deployed in low cost embedded platforms with limited computing resources.
@@
Many approaches have been proposed to deal with the computational cost of MPC and facilitate its use in real-world applications. These approaches can be broadly categorised as efficient optimisation solvers, Explicit MPC (EMPC), and approximate optimisation methods. Efficient solvers are specialised optimisation routines designed for specific problem formulations; they generally make use of hot start techniques and early stopping to reduce computation times. This type of approach can succeed in enabling some MPC implementations in real time but for complex forms of MPC or where long horizons are considered, the reduction in computation time is often insufficient. For more information on such solvers, refer to the review in Ferreau et al. (2017). Alternatively, EMPC is a form of MPC where the optimisation problem is solved offline and stored as lookup tables that map system states and target outputs to optimal control inputs. This can be done exactly for Linear MPC (LMPC)  Bemporad et al. (2002) and approximately for NMPC  Johansen (2004); Bemporad and Filippi (2006). However, producing such lookup tables requires solving large multi-parametric programs which often become intractable for large systems, and when solutions are possible, lookup tables grow exponentially with horizon length to sizes that do not fit in the limited memory of embedded controllers, or that are still too slow to query in real time Karg and Lucia (2020a); Fabiani and Goulart (2022). For a survey of EMPC refer to Alessio and Bemporad (2009). Lastly, approximate optimisation methods are approaches where conventional solvers are replaced with alternatives that can produce similar outputs for the same inputs but are much cheaper to evaluate in terms of computation cost. A particularly effective way of doing this is to replace the optimisation solver by a Neural Network (NN) that has been trained to imitate the solver thereby reducing the optimisation operation to a simple function evaluation, Figure 1 illustrates how this idea can be applied to MPC. We refer to this type of approach as Neural Network-Based Optimisation (NNBO), neural solvers, or neural approximations.
@@
The use, benefits, and limitations of efficient optimisation solvers and EMPC have been presented and discussed on previous reviews Ferreau et al. (2017); Alessio and Bemporad (2009), yet a review on the use of NNs as a replacement of the solver in MPC remains absent in scientific literature. Norouzi et al. Norouzi et al. (2023) touched on this topic in their review of ML in MPC for automotive applications. However, the discussion on the matter was limited to four studies which do not encompass all of the existing approximation frameworks. Similarly, Hewing et al. Hewing et al. (2020) and Mesbah et al. Mesbah et al. (2022) briefly discussed the subject in their reviews on ways in which ML has been integrated in MPC. But again, the scope of those works is broad and does not go into detail on the subject of NNBO. Herein, a survey is presented of the various frameworks and methods that can be used to generate approximate MPC controllers powered by NNBO. Compared to the aforementioned works Norouzi et al. (2023); Hewing et al. (2020); Mesbah et al. (2022), we focus exclusively on this topic. Moreover, we present and compare a wide range of unique existing frameworks including discussions on how the different approaches deal with the challenges of NNBO. Our contributions can be summarised as follows:
@@
The remainder of the paper is organised as follows, Section 2 explains the scope of the survey and the type of method that it covers with reference to related surveys that cover out-of-scope topics. Section 3 provides a general introduction to MPC and introduces our proposed taxonomy of NNBO methods. In Sections 4 – 6, we introduce and explain different frameworks that can be used to approximate LMPC, NMPC, and RMPC with NNBO, respectively. We compare the different methods according to the type of guarantees they offer, the size of problem they have been validated on, the factor by which they speed up the controller, and evidence showcasing their success beyond simulation.
@@ 2 Scope and Related Work
Machine Learning (ML) techniques have been extensively used in the field of MPC for various purposes besides accelerating optimisation, therefore, the scope of this survey excludes certain topics. First of all, we differentiate between approaches where the MPC optimiser is replaced with an NN that requires a single evaluation (e.g. Feedforward NNs), and approaches stemming from the work by Tank and Hopfield Tank and Hopfield (1986), where the optimiser is replaced by a Recurrent NN (RNN) that needs to be simulated iteratively until convergence to yield a prediction. This survey focuses on approaches that necessitate a single evaluation. Exhaustive surveys on RNNs for optimisation, often referred to as Hopfield NNs, can be found in references Wen et al. (2009); Jin et al. (2019). Furthermore, we omit publications that utilise ML to construct data-driven models for use in MPC, unless these methods are integrated within a framework that concurrently approximates and expedites the optimiser. Although data-driven models can lead to reductions in computation time, their purpose is to achieve better plant predictive accuracy, not sufficient speed for implementation in real time. The works in Ren et al. (2022); Norouzi et al. (2023); Mesbah et al. (2022); Hewing et al. (2020); Meng et al. (2022) provide good coverage of such approaches. In addition, we do not cover the topic of using ML techniques, or most commonly Reinforcement Learning (RL), to find parametrisations of the cost function in MPC that yield better performance or enable online controller adaptation. Approaches of this nature have been assessed in previous reviews Mesbah et al. (2022); Hewing et al. (2020) and are designed to enhance controllers that are already suitable for real-time deployment, rather than facilitating their initial deployment. However, where relevant we do discuss the subject of improving neural approximations of MPC with RL. Similarly, we only briefly touch on the topic of initialising RL agents with behaviours learnt from MPC Karg and Lucia (2021). Such approaches are an extension of the methods reviewed here and a complete review is outside the scope of the paper. We also note that there are several publications (e.g. Chakrabarty et al. (2016); Bemporad et al. (2011); Csekő et al. (2015); Karg and Lucia (2020a)) on the topic of learning EMPC policies to reduce the computational burden of the lookup operations at run time. We exclude such approaches from our scope because they rely on being able to generate an EMPC controller in the first place, which is not always possible. The methods discussed here are not constrained by this limitation. Lastly, we only provide a high level introduction to MPC and the different variants that exist. For an in depth dive into the theory of MPC, readers are directed to reference Kouvaritakis and Cannon (2015). Those seeking an application-oriented perspective are directed to references Schwenzer et al. (2021); Rakovic and Levine (2018).
@@ 3.1 Model Predictive Control Basics
An MPC controller implementation typically requires three components: an optimisation solver, a plant model, and a state estimator, as shown in the block diagram in Figure 2. The optimisation solver is one of many existing algorithms chosen according to the type of problem involved and it is used to solve the underlying optimisation problem in MPC. Commonly encountered problem types in MPC are Quadratic Programs (QPs), Nonlinear Programming (NLP) problems, min-max problems, and Mixed Integer Linear Programs (MILPs)  Ferreau et al. (2017); Kouvaritakis and Cannon (2015). The plant model is a function that captures the dynamics of the system being controlled and enables the simulation of the system responses to inputs. Lastly, the state estimator is an algorithm that can be used to approximate system states and disturbances that cannot be directly measured from other variables that can be measured. The most popular realisation of this subsystem is a Kalman filter Rakovic and Levine (2018) but it should be noted that a state estimator is not always required; in some systems all variables are directly measurable. From hereon, we focus the discussion on the optimiser and plant blocks.
@@
The generalised form of a discrete-time plant model is given by (3.1), where the subscript $i$ indicates the current time step, $t$ is the current time, and $x\in\mathbb{R}^{n_{x}}$, $u\in\mathbb{R}^{n_{u}}$, and $y\in\mathbb{R}^{n_{y}}$ are the state, input and output vectors of the system Hewing et al. (2020). The uncertainty of the system is characterised by the variables $\theta$, $w$, and $v$. More specifically, $\theta$ accounts for parametric model uncertainties and is constant over time, $w$ accounts for disturbances and process noise, and $v$ accounts for output measurement noise. Both $w$ and $v$ are made up of independent and identically distributed (i.i.d). time varying random variables. Lastly, the functions $f$ and $g$ are the state and output transition functions, which may take any arbitrary form, for instance a linear state space model, a neural network, a lookup table, etc.
$x_{i+1}=f(x_{i},u_{i},t,w_{i},\theta)$ (3.1)
@@
At each control interval the optimiser uses the plant model to solve an optimisation problem of the form shown in (3.1)
$\min_{U}\quad$ $C(U)=\sum_{i=0}^{N_{p}-1}l(x_{i},u_{i},i)+l_{f}(x_{N_{P}},u_{N_{P}},N_{p})$ (3.1)
@@
where $l$ and $l_{f}$ are the stage and terminal cost functions, $N_{p}$ is the prediction horizon length (i.e. the number of time steps that the controller looks forward into the future), $\mathcal{U}_{j}$, $\mathcal{Y}_{j}$ and $\mathcal{X}_{j}$ are generalised forms of equality and inequality constraints expressed as sets, and $\mathcal{X}_{f}$ is the terminal states set. The cost function $l$ is typically designed to achieve control objectives such as reference tracking, trajectory tracking, path tracking, control input smoothing, and control input minimisation. For each state, output, and input one may choose to use none, one, or several of these objectives depending on the desired behaviour. The equality and inequality constraints (2a - 2e) are used to enforce the system dynamics and all input, output and state constraints over the horizon. Lastly, the constraint  (2f) and the cost $l_{f}$ are used to enforce terminal conditions, which, in certain types of MPC, provide assurances of stability and recursive feasibility. The solution of (3.1) is the optimal sequence of control inputs ($U^{*}$) to be applied to the system which will drive it to its desired states over the prediction horizon. In practice, only the first element of $U^{*}$ is applied to the system and the entire control sequence is recomputed at the next control interval by solving the optimisation problem again. It should be noted that many MPC implementations also distinguish between the prediction horizon and the control horizon, which is a shorter window of future time of length ($N_{c}$) over which control actions are optimised, that is ($N_{c}<N_{p}$akovic and Levine (2018).
In the following subsections, we provide a short description of the different types of MPC, including detail about the specific forms of (3.1) and (3.1) that are used. This information is important to understand the benefits and challenges of replacing the optimiser block in Figure 2 with an NNBO method.
@@ 3.1.1 Linear Model Predictive Control
Linear MPC (LMPC) is the most basic form of MPC. It is limited to employing linear plant models alongside a quadratic cost function coupled with linear constraints. The typical choice of plant is a time invariant state space model of the form shown in Equation (3Kouvaritakis and Cannon (2015). However, it is also feasible to utilise time-varying plants. For time invariant plants the matrices $A$, $B$, $C$, and $D$ in (3) are fixed, whereas for time-varying plants these matrices can change over time (e.g. $A=A(t)$). In practice, time-varying plants can be used to linearise complex systems at various operating points, thus offering more accurate modelling. Due to this distinction, we refer to LMPC that uses a time-varying plant as Linear Time Varying MPC (LVMPC).
@@
Through algebraic manipulation Equation (4) can be restructured into a standard QP the form of which is shown in (5). For LMPC $P$, $q$, and $R$, are fixed. Whereas, for LVMPC, these matrices vary over time to capture the changing dynamics of the plant. The format in (5) enables much more efficient optimisation and many optimisers can solve large versions of this problem in real time at high frequencies, especially if the problem is strictly convex Ferreau et al. (2014); Stellato et al. (2020).
@@ 3.1.2 Nonlinear Model Predictive Control
NMPC is a more flexible formulation of MPC that allows the use of nonlinear plant models, any form of cost function, and nonlinear constraints Alessio and Bemporad (2009). The plant model need not conform to a specific structure (e.g. a state space model), it can be represented in any manner that is most suitable or efficient for addressing the problem at hand. The generalised form of an NMPC plant and optimisation problem is already well captured by Equations (3.1) and (3.1). The higher flexibility of NMPC comes at the cost of increased difficulty in solving the underlying optimisation problem which is typically of the NLP type. In theory, the plant, cost and constraints can take any arbitrary form, but in practice most implementations restrict these choices to forms that are cheap to evaluate and result in problems that can be solved efficiently, for instance using Sequential Quadratic Programming (SQP). Thus, the capabilities of NMPC are rarely fully exploited in real-world applications, NNBO can help address this limitation.

### v3

@@ Abstract
[…]
@@ 1 Introduction
Model Predictive Control (MPC) is an optimal control scheme that is widely used in the process industry and robotics. Compared to other control schemes, its main advantages include seamless handling of MIMO systems, ease of consideration of state, input and output constraints, the guarantee of closed loop stability by design, and the ability to exploit knowledge and predictions of future control references or disturbances (Schwenzer et al., 2021; Kouvaritakis and Cannon, 2015). In a nutshell, MPC controllers use a model or “plant” of the controlled system to derive a sequence of control inputs that will result in optimal system response over a future period of time, or “horizon”, given the current states of the system and desired reference trajectories. This is achieved by solving an optimisation problem where the design variables are the sequence of future system inputs and the plant is used to calculate future system outputs and states for said sequence. The primary challenge in deploying MPC controllers in practical applications stems from the necessity to solve large optimisation problems in real time. For systems with fast dynamics requiring control at high frequency, this control scheme becomes infeasible even for simple systems, especially when the more complex formulations of MPC such as Nonlinear MPC (NMPC) or Robust MPC (RMPC) are used, or when the controller needs to be deployed in low cost embedded platforms with limited computing resources.
@@
Many approaches have been proposed to deal with the computational cost of MPC and facilitate its use in real-world applications. These approaches can be broadly categorised as efficient optimisation solvers, Explicit MPC (EMPC), and approximate optimisation methods. Efficient solvers are specialised optimisation routines designed for specific problem formulations; they generally make use of hot start techniques and early stopping to reduce computation times. This type of approach can succeed in enabling some MPC implementations in real time but for complex forms of MPC or where long horizons are considered, the reduction in computation time is often insufficient. For more information on such solvers, refer to the review by Ferreau et al. (2017). Alternatively, EMPC is a form of MPC where the optimisation problem is solved offline and stored as lookup tables that map system states and target outputs to optimal control inputs. This can be done exactly for Linear MPC (LMPC) (Bemporad et al., 2002) and approximately for NMPC (Johansen, 2004; Bemporad and Filippi, 2006). However, producing such lookup tables requires solving large multi-parametric programs which often become intractable for large systems, and when solutions are possible, lookup tables grow exponentially with horizon length to sizes that do not fit in the limited memory of embedded controllers, or that are still too slow to query in real time (Karg and Lucia, 2020a; Fabiani and Goulart, 2022). For a survey of EMPC refer to Alessio and Bemporad (2009). Lastly, approximate optimisation methods are approaches where conventional solvers are replaced with alternatives that can produce similar outputs for the same inputs but are much cheaper to evaluate in terms of computation cost. A particularly effective way of doing this is to replace the optimisation solver by a Neural Network (NN) that has been trained to imitate the solver thereby reducing the optimisation operation to a simple function evaluation, Figure 1 illustrates how this idea can be applied to MPC. We refer to this type of approach as Neural Network-Based Optimisation (NNBO), neural solvers, or neural approximations.
@@
The use, benefits, and limitations of efficient optimisation solvers and EMPC have been presented and discussed on previous reviews (Ferreau et al., 2017; Alessio and Bemporad, 2009), yet a review on the use of NNs as a replacement of the solver in MPC remains absent in literature. Norouzi et al. (2023) touched on this topic in their review of ML in MPC for automotive applications. However, the discussion on the matter was limited to four studies which do not encompass all of the existing approximation frameworks. Similarly, Hewing et al. (2020) and Mesbah et al. (2022) briefly discussed the subject in their reviews on ways in which ML has been integrated in MPC. But again, the scope of those works is broad and does not go into detail on the subject of NNBO. Herein, a survey is presented of the various frameworks and methods that can be used to generate approximate MPC controllers powered by NNBO. Compared to the aforementioned works (Norouzi et al., 2023; Hewing et al., 2020; Mesbah et al., 2022), we focus exclusively on this topic. Moreover, we present and compare a wide range of unique existing frameworks including discussions on how the different approaches deal with the challenges of NNBO. Our contributions can be summarised as follows:
@@
The remainder of the paper is organised as follows, Section 2 explains the scope of the survey and the type of method that it covers with reference to related surveys that cover out-of-scope topics. Section 3 provides a general introduction to MPC and introduces our proposed taxonomy of NNBO methods. In Sections 46, we introduce and explain different frameworks that can be used to approximate LMPC, NMPC, and RMPC with NNBO, respectively. We compare the different methods according to the type of guarantees they offer, the size of problem they have been validated on, the factor by which they speed up the controller, and evidence showcasing their success beyond simulation.
@@ 2 Scope and Related Work
Machine Learning (ML) techniques have been extensively used in the field of MPC for various purposes besides accelerating optimisation, therefore, the scope of this survey excludes certain topics. First of all, we differentiate between approaches where the MPC optimiser is replaced with an NN that requires a single evaluation (e.g. Feedforward NNs), and approaches stemming from the work by Tank and Hopfield (1986), where the optimiser is replaced by a Recurrent NN (RNN) that needs to be simulated iteratively until convergence to yield a prediction. This survey focuses on approaches that necessitate a single evaluation. Exhaustive surveys on RNNs for optimisation, often referred to as Hopfield NNs, have been presented by Wen et al. (2009) and Jin et al. (2019). Furthermore, we omit publications that utilise ML to construct data-driven models for use in MPC, unless these methods are integrated within a framework that concurrently approximates and expedites the optimiser. Although data-driven models can lead to reductions in computation time, their purpose is to achieve better plant predictive accuracy, not sufficient speed for implementation in real time. The works by Ren et al. (2022); Norouzi et al. (2023); Mesbah et al. (2022); Hewing et al. (2020) and Meng et al. (2022) provide good coverage of such approaches. In addition, we do not cover the topic of using ML techniques, or most commonly Reinforcement Learning (RL), to find parametrisations of the cost function in MPC that yield better performance or enable online controller adaptation. Approaches of this nature have been assessed in previous reviews (Mesbah et al., 2022; Hewing et al., 2020) and are designed to enhance controllers that are already suitable for real-time deployment, rather than facilitating their initial deployment. However, where relevant we do discuss the subject of improving neural approximations of MPC with RL. Similarly, we only briefly touch on the topic of initialising RL agents with behaviours learnt from MPC (Karg and Lucia, 2021). Such approaches are an extension of the methods reviewed here and a complete review is outside the scope of the paper. We also note that there are several publications (e.g. (Chakrabarty et al., 2016; Bemporad et al., 2011; Csekő et al., 2015; Karg and Lucia, 2020a)) on the topic of learning EMPC policies to reduce the computational burden of the lookup operations at run time. We exclude such approaches from our scope because they rely on being able to generate an EMPC controller in the first place, which is not always possible. The methods discussed here are not constrained by this limitation. Lastly, we only provide a high level introduction to MPC and the different variants that exist. For an in depth dive into the theory of MPC, readers are directed (Kouvaritakis and Cannon, 2015). Those seeking an application-oriented perspective are directed to Schwenzer et al. (2021) and Rakovic and Levine (2018).
@@ 3.1 Model Predictive Control Basics
An MPC controller implementation typically requires three components: an optimisation solver, a plant model, and a state estimator, as shown in the block diagram in Figure 2. The optimisation solver is one of many existing algorithms chosen according to the type of problem involved and it is used to solve the underlying optimisation problem in MPC. Commonly encountered problem types in MPC are Quadratic Programs (QPs), Nonlinear Programming (NLP) problems, min-max problems, and Mixed Integer Linear Programs (MILPs) (Ferreau et al., 2017; Kouvaritakis and Cannon, 2015). The plant model is a function that captures the dynamics of the system being controlled and enables the simulation of the system responses to inputs. Lastly, the state estimator is an algorithm that can be used to approximate system states and disturbances that cannot be directly measured from other variables that can be measured. The most popular realisation of this subsystem is a Kalman filter (Rakovic and Levine, 2018) but it should be noted that a state estimator is not always required; in some systems all variables are directly measurable. From hereon, we focus the discussion on the optimiser and plant blocks.
@@
The generalised form of a discrete-time plant model is given by (1), where the subscript $i$ indicates the current time step, $t$ is the current time, and $x\in\mathbb{R}^{n_{x}}$, $u\in\mathbb{R}^{n_{u}}$, and $y\in\mathbb{R}^{n_{y}}$ are the state, input and output vectors of the system (Hewing et al., 2020). The uncertainty of the system is characterised by the variables $\theta$, $w$, and $v$. More specifically, $\theta$ accounts for parametric model uncertainties and is constant over time, $w$ accounts for disturbances and process noise, and $v$ accounts for output measurement noise. Both $w$ and $v$ are made up of independent and identically distributed (i.i.d). time varying random variables. Lastly, the functions $f$ and $g$ are the state and output transition functions, which may take any arbitrary form, for instance a linear state space model, a neural network, a lookup table, etc.
$x_{i+1}=f(x_{i},u_{i},t,w_{i},\theta)$ (1)
@@
At each control interval the optimiser uses the plant model to solve an optimisation problem of the form shown in (2)
$\min_{U}\quad$ $C(U)=\sum_{i=0}^{N_{p}-1}l(x_{i},u_{i},i)+l_{f}(x_{N_{P}},u_{N_{P}},N_{p})$ (2)
@@
where $l$ and $l_{f}$ are the stage and terminal cost functions, $N_{p}$ is the prediction horizon length (i.e. the number of time steps that the controller looks forward into the future), $\mathcal{U}_{j}$, $\mathcal{Y}_{j}$ and $\mathcal{X}_{j}$ are generalised forms of equality and inequality constraints expressed as sets, and $\mathcal{X}_{f}$ is the terminal states set. The cost function $l$ is typically designed to achieve control objectives such as reference tracking, trajectory tracking, path tracking, control input smoothing, and control input minimisation. For each state, output, and input one may choose to use none, one, or several of these objectives depending on the desired behaviour. The equality and inequality constraints (2a) - (2e) are used to enforce the system dynamics and all input, output and state constraints over the horizon. Lastly, the constraint (2f) and the cost $l_{f}$ are used to enforce terminal conditions, which, in certain types of MPC, provide assurances of stability and recursive feasibility. The solution of (2) is the optimal sequence of control inputs ($U^{*}$) to be applied to the system which will drive it to its desired states over the prediction horizon. In practice, only the first element of $U^{*}$ is applied to the system and the entire control sequence is recomputed at the next control interval by solving the optimisation problem again. It should be noted that many MPC implementations also distinguish between the prediction horizon and the control horizon, which is a shorter window of future time of length ($N_{c}$) over which control actions are optimised, that is ($N_{c}<N_{p}$) (Rakovic and Levine, 2018).
In the following subsections, we provide a short description of the different types of MPC, including detail about the specific forms of (1) and (2) that are used. This information is important to understand the benefits and challenges of replacing the optimiser block in Figure 2 with an NNBO method.
@@ 3.1.1 Linear Model Predictive Control
Linear MPC (LMPC) is the most basic form of MPC. It is limited to employing linear plant models alongside a quadratic cost function coupled with linear constraints. The typical choice of plant is a time invariant state space model of the form shown in (3) (Kouvaritakis and Cannon, 2015). However, it is also feasible to utilise time-varying plants. For time invariant plants the matrices $A$, $B$, $C$, and $D$ in (3) are fixed, whereas for time-varying plants these matrices can change over time (e.g. $A=A(t)$). In practice, time-varying plants can be used to linearise complex systems at various operating points, thus offering more accurate modelling. Due to this distinction, we refer to LMPC that uses a time-varying plant as Linear Time Varying MPC (LVMPC).
@@
Through algebraic manipulation (4) can be restructured into a standard QP the form of which is shown in (5). For LMPC $P$, $q$, and $R$, are fixed. Whereas, for LVMPC, these matrices vary over time to capture the changing dynamics of the plant. The format in (5) enables much more efficient optimisation and many optimisers can solve large versions of this problem in real time at high frequencies, especially if the problem is strictly convex (Ferreau et al., 2014; Stellato et al., 2020).
@@ 3.1.2 Nonlinear Model Predictive Control
NMPC is a more flexible formulation of MPC that allows the use of nonlinear plant models, any form of cost function, and nonlinear constraints (Alessio and Bemporad, 2009). The plant model need not conform to a specific structure (e.g. a state space model), it can be represented in any manner that is most suitable or efficient for addressing the problem at hand. The generalised form of an NMPC plant and optimisation problem is already well captured by Equations (1) and (2). The higher flexibility of NMPC comes at the cost of increased difficulty in solving the underlying optimisation problem which is typically of the NLP type. In theory, the plant, cost and constraints can take any arbitrary form, but in practice most implementations restrict these choices to forms that are cheap to evaluate and result in problems that can be solved efficiently, for instance using Sequential Quadratic Programming (SQP). Thus, the capabilities of NMPC are rarely fully exploited in real-world applications, NNBO can help address this limitation.